Circular No. 7979

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002fs, 2002ft, AND 2002fu Further to IAUC 7905, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of three apparent supernova on unfiltered Palomar NEAT images, as tabulated below; K. Li was also involved with SN 2002ft.

SN  2002 UT  α_{2000}  δ_{2000}  Mag.  Offset
2002fs Sept. 3.26 22^{h}54^{m}00.^{s}88 +14^{o}39′24′′.5 20.1 2′′.5 E
2002ft  9.39 23 15 45.82 −16 55 35.2 18.9 1′′ E, 2′′ S
2002fu  10.21 22 09 55.99 +12 05 06.6 19.4 5′′ W, 1′′ N

The offset of SN 2002ft is with respect to “a faint nearby object”, as it is unclear whether the presumed galaxy center is actually a galaxy or two blended stars. Additional magnitudes from NEAT images: SN 2002fs, 2001 Oct. 6, [20.5; 2002 Aug. 4, [20.5; Sept. 11, 17.4. SN 2002ft, Aug. 30, [20.5; Sept. 17, 18.1. SN 2002fu, 2001 Nov. 5, [20.5; 2002 Aug. 22, [20.5; Sept. 3, [20.5; 11, 19.3.

M. Hamuy, Carnegie Observatories, reports that spectra (range 360– 900 nm) of these new objects were obtained on Sept. 28.18–28.23 UT with the Baade 6.5-m telescope (+ LDSS-2) at Las Campanas. SN 2002fu is a type-II supernova with a broad Hα line exhibiting a P-Cyg profile; the minimum of the absorption component yields an expansion velocity of 14000 km/s. A narrow Hα emission from the host galaxy yields a redshift of 0.091. SN 2002fs bears resemblance to the type-Ia SN 1994D, five days past maximum; a narrow Hα emission from the host galaxy yields a redshift of 0.038. SN 2002ft also bears resemblance to SN 1994D, two days past maximum; by comparison to SN 1994D, a redshift of ≈ 0.075 is derived.

SUPERNOVAE 2002fi AND 2002fj Hamuy also writes that spectra were obtained, as above, for these

two supernovae on Sept. 26.29 and 26.39 UT. SN 2002fi (cf. IAUC 7972) is a type-Ia supernova whose spectrum bears strong resemblance to SN 1994D, 20 days past maximum. SN 2002fj (cf. IAUC 7973) is of type IIn, its spectrum dominated by a strong Hα emission line of intermediate (700 km/s) width without appreciable Doppler shift relative to the host galaxy; this emission is superposed on a broad component. No signs of He I lines are present.

2002 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green